OpenStudy (anonymous):

an elevator car whose floor to ceiling distance is equal to 2.7 m , starts from rest and ascends with a constant acceleration of 1.2 ms^2 .Two seconds later , a bolt drops from the car's ceiling .The time taken by the bolt to hit the floor of the elevator is a. 1 s b. 0.7 s

OpenStudy (anonymous):

The actual speed at the start doesn't matter, since the bolts relative speed to the car is zero, when it gets loose. So the cars acceleration + the gravitational acceleration gives the bolt's acceleration relative to the car: \[a = g+ 1.2 m /s ^{2} = 11.2 m/s^2\] \[s = a*t^2/2\] \[t=\sqrt{2s/a}= 0.69\approx0.7\] therefore, it's b

OpenStudy (anonymous):

thank u.. :)